# Dedelayed: Deleting remote inference delay via on-device correction

### v1

@@ Abstract
Remote inference allows lightweight devices to leverage powerful cloud models. However, communication network latency makes predictions stale and unsuitable for real-time tasks. To address this, we introduce Dedelayed, a delay-corrective method that mitigates arbitrary remote inference delays, allowing the local device to produce low-latency outputs in real time. Our method employs a lightweight local model that processes the current frame and fuses in features that a heavyweight remote model computes from past frames. On video from the BDD100K driving dataset, Dedelayed improves semantic segmentation accuracy over the stronger of the local-only and remote-only baselines across all realistic communication network delays beyond 33 ms. Without incurring additional delay, it improves accuracy by 6.4 mIoU compared to fully local inference and 9.8 mIoU compared to remote inference, for a round-trip delay of 100 ms. The advantage grows under longer delays and higher-motion scenes, as delay-mitigated split inference sustains accuracy more effectively, providing clear advantages for real-time tasks that must remain aligned with the current world state.
@@ 1 Introduction
In soft real-time applications—such as cloud gaming or video conferencing—late outputs may be diminished in value, but are still useful. In these applications, we can offload expensive computations to powerful cloud GPUs to save on-device power. As long as the typical latency is low, the loss of utility from latency is outweighed by power savings and extended battery life. However, in hard real-time applications—such as aerial robotic control or obstacle avoidance—late outputs can be catastrophic, and the system must be designed with a guaranteed deadline. Due to the irreducible high-tail latency in wireless communication, hard real-time applications must be equipped with a fully functional local inference pipeline as a fallback in the case that the remote predictions fail to meet the deadline. In this work, we focus on such real-time applications running on resource-constrained devices, relying on inference using video inputs.
In recent years, various approaches to split computing have been proposed to offload computation of expensive image and video models to the cloud to enable the next-generation of robotic, remote sensing, and wearable technology platforms. For real-time streaming video applications, existing approaches still fall into one or more of three common pitfalls. (1) They allocate all on-device power and computation to a single linear inference pipeline, leaving no resources for a local-only fallback. (2) They do not account for the impact of latency on prediction accuracy. (3) They operate on videos with significantly reduced spatiotemporal resolution to manage computational cost, leaving out rich visual details available from modern camera systems.
Figure 1: Overview of the Dedelayed real-time inference setup. The lightweight local model and a powerful remote model augment each other’s strengths to produce accurate and timely outputs.
To address these limitations, we introduce Dedelayed (Fig.˜1), a co-inference framework that leverages fresh local information to mitigate the effects of remote inference delay. It consists of a local model and a remote model connected over a communication network, and fuses delayed remote signals into a local pipeline operating on fresh on-device inputs, yielding real-time performance that is never worse than either model in isolation. This enables the use of high-capacity cloud models for delay-sensitive applications while avoiding the pitfalls of prior offloading approaches:
@@
    Temporal prediction for latency mitigation. During supervised training of the remote model component, we simulate a delay of $D$ frames. In other words, the remote model is trained to predict the future. A delay embedding—similar to a position embedding in text or vision transformers—allows the behavior of the remote model to adapt to changes in the channel. As shown in Fig.˜2, temporally predictive training is able to capture motion dynamics, which can be used to compensate for latency.
@@
    Mixed-resolution inference. On-device AI video processing at or near the capture resolution and frame rate—typically >1>1 megapixel and >20>20 frames per second—is rarely feasible, even with lightweight models. To save resources, real-time computer vision applications often process each frame independently or use small motion updates instead of natively processing a dense 3D pixel volume. Dedelayed is capable of using a mixed-resolution—instead of reducing the resolution for the entire inference pipeline, only the local model resolution is reduced. In parallel, the remote model operates on multiple, high-resolution frames using a 3D transformer. Thus, the remote model can utilize powerful GPUs to model fine details and motion of delayed frames, while the local component can allocate its resources to modeling the current state of objects and the scene, as shown in Fig.˜3.
@@
    We introduce Dedelayed, a split computing framework for video inputs that integrates the output of a future-predicting remote model with the current output of a local model.
@@
For experimental validation, we demonstrate a simple Dedelayed system with an addition-based fusion on off-the-shelf models with minimal architectural changes. This makes it easy to enhance existing pipelines, deploy in practice, and extend to other real-time methods.
@@ 2 Background
In the human visual system, the optic nerve can only transmit a small fraction of the information received by the retina (Kelly, 1962). Barlow’s efficient coding hypothesis (Barlow et al., 1961) posits that compression is the primary role of early processing; once this compressed representation is received in deeper layers of the visual cortex, more metabolically intense processing can occur. In the predictive coding model (Rao & Ballard, 1999), this processing is driven by feedback mechanisms that minimize a temporally predictive error signal to create a perceptual model that is consistent with sensory inputs.
Machines equipped with digital video sensors—which are at the center of ongoing innovation in robotics (Kim et al., 2024; O’Neill et al., 2024), remote sensing (Szwarcman et al., 2024; Khani et al., 2021), and wearable technology (Grauman et al., 2022; 2024)—share similar constraints. The throughput and power efficiency of ingesting pixels on the sensor device (e.g., a battery-powered robot) are extremely high—typically tens or hundreds of megapixels per watt-second (Engel et al., 2023). However, moderately sized DNNs can only process visual data at about one megapixel per watt-second (Cai et al., 2023). For more advanced video AI based on autoregressive modeling (Agarwal et al., 2025) or temporal prediction (Assran et al., 2025), the efficiency may be as low as 500 pixels per watt-second. Instead of on-device processing, power constraints can be circumvented by compressing and transmitting video streams to cloud GPU datacenters supported by a 100-megawatt power infrastructure (Goldberg & Kehoe, 2013).
However, fully remote processing is challenging for certain real-time applications (e.g., collision avoidance) due to unreliability in network and cloud infrastructure (Chen et al., 2024; 2025). Thus, delivering predictions by a guaranteed deadline requires a fallback procedure independent of the remote server. In many systems (e.g., autonomous motor vehicles) the limited accuracy and reliability of lightweight local models warrant a human operator as the fallback (Committee, 2021), preventing full automation.
@@ 3 Related work
Prior work has extensively explored lightweight architectures (e.g., EfficientViT (Cai et al., 2023), MobileNetV4 (Qin et al., 2024)) that minimize computation to achieve real-time on-device performance. These approaches deliver low latency but are constrained by device power and compute. When devices are too limited, the common alternative is fully remote inference, which offloads computation to servers, but is highly susceptible to network latency.
Split-computing approaches largely focus on distributing workloads rather than optimizing for strict real-time operation. Next-generation compression standards such as MPEG AI (ISO/IEC, 2025) and JPEG AI (Ascenso et al., 2023) target bandwidth reduction via task-specific compression, significantly lowering transmission costs with reasonable compute overhead. However, even they lack explicit mechanisms to anticipate or compensate for network delay, leading to stale predictions misaligned with the current world state.
Other efforts related to our work have also been explored. Clockwork Convnets (Shelhamer et al., 2016) reuse stale features to reduce inference latency, but they offer limited temporal reasoning and operate on a single device. Accel (Jain et al., 2019) warps heavy-model features forward with optical flow and corrects them with a lightweight model, but is also not intended for across-network operation. Adaptive Model Streaming (Khani et al., 2021) streams weight updates from a server to keep a local model fresh, focusing on model adaptation rather than directly mitigating per-frame staleness from communication latency. Though less known, Knowledge Boosting (Srinivas et al., 2024) is very closely related to our work. Like us, it fuses delayed remote features with a small on-device model, but it assumes a fixed delay. We generalize to longer and variable latencies by conditioning on a tunable delay while keeping the design simple and reusable.
@@ 4.1 Delay mitigation framework
Dedelayed introduces a general framework that improves the accuracy and robustness of real-time inference on resource-constrained sensor devices. It does so by combining the strengths of both local inference and remote inference, while mitigating their weaknesses. The local model has access to the latest sensor data, and yet lacks the computational capability needed to produce accurate outputs. The remote model provides accurate outputs, and yet delivers them with delay. By careful combination of both subsystems, Dedelayed is able to provide bounded performance guarantees—it is never worse than either local inference or remote inference independently. As we will demonstrate later, we are able to glue together the two subsystems in a way that is simple yet effective.
@@
Dedelayed can be formulated in simple mathematical terms as follows. Given a fresh input frame $x_{t}$ at current time $t$, the final prediction $\hat{y}_{t}$ is computed using a lightweight local model, $f_{\text{light}}$, which processes $x_{t}$ along with time-delayed features $z_{t-\tau}$ from a heavyweight remote model, $f_{\text{heavy}}$. To produce powerful predictive features, the remote model is conditioned on the delay $\tau$, and processes a short clip of past frames $x_{\leq t-\tau}$ ending at time $t-\tau$. This is expressed by the following equations:
$z_{t-\tau}=f_{\text{heavy}}(\tau,\,x_{\leq t-\tau})$ (1)
$\hat{y}_{t}=f_{\text{light}}(x_{t},\,z_{t-\tau})$ (2)
For clarity, the notation is summarized in Table˜1. Fig.˜4 presents a system diagram that demonstrates the fundamental principle we describe, and shows how information propagates through the various subsystems as time progresses.
@@
$z_{t-\tau}$ Features outputted by heavyweight model run at time $t-\tau$
$\tau$ Delay in time between the old and current frame at time $t$
$D$ Delay in frames between the old and current frame at time $t$
$f_{\text{light}}$ Lightweight model run at time $t$
$f_{\text{heavy}}$ Heavyweight model run at time $t-\tau$
@@
4.2 Design and Implementation
Dedelayed consists of a lightweight on-device model and a predictive remote model connected over a communication network. The remote side computes delay-conditioned features from past inputs and returns them to the device, where they are fused early into the local model running on the current input. To provide a guarantee on baseline performance, these components can first be trained to work independently and later fused and trained jointly.
@@ 4.2 Design and Implementation
    Each incoming frame is independently encoded into features using a pretrained 2D ViT backbone on the remote. We maintain a context window of the $K$ most recent features.
@@
    The lightweight local model runs on a fresh input $x_{t}$, and fuses in the remote features $z_{t-\tau}$.
@@
Remote predictive module.
Figure Fig.˜5 visualizes the remote component. The remote model processes a fixed context of $K$ past frames ending at $t-\tau$. Each frame is encoded independently with a 2D ViT—we use EfficientViT-L1 with an effective $8\times 8$ patch. The per-frame features are concatenated along the temporal axis and spatially merged into larger $16\times 16$ patches to keep the sequence length comparable. (When $K=4$, the sequence length is identical.) A learned delay embedding determined by the delay $\tau$ is added, and the result is then processed by a 3D ViT followed by learned pooling (MLP–pool–MLP) to produce delay-conditioned features $z_{t-\tau}$. We pretrain the remote predictive module by attaching a task head to the 3D ViT backbone and training it to predict the target labels $y_{t}$ from inputs up to time $t-\tau$.
Figure 5: Overview of the remote video predictive model, trained to predict the label at index $n$ from a $K$-frame context ending at nDn-D.
Local model and fusion.
The lightweight local model processes the fresh input $x_{t}$. We compute first-stage features $h=\texttt{T1}(x_{t})$ and perform early fusion by element-wise addition with the delayed remote features, $h^{\prime}=h+z_{t-\tau}$. Both tensors have shape $96\times H/8\times W/8$, so no projection or resizing is required. The fused features $h^{\prime}$ are then processed by the remaining local blocks and decoded to the final output $\hat{y}_{t}$. If $z_{t-\tau}$ is unavailable, the local model falls back to $h^{\prime}=h$ and proceeds unchanged.
5 Experiments
We focus our evaluation of Dedelayed on the task of real-time semantic segmentation of driving scenes, a domain where timely and accurate perception is critical for applications such as autonomous driving and robotics. We demonstrate that delay-aware feature fusion can mitigate remote inference latency, sustaining accuracy even when remote predictions are delayed by long communication network latencies.

### v2

@@ Abstract
Video comprises the vast majority of bits that are generated daily, and is the primary signal driving current innovations in robotics, remote sensing, and wearable technology. Yet, the most powerful video understanding models are too expensive for the resource-constrained platforms used in these applications. One approach is to offload inference to the cloud; this gives access to GPUs capable of processing high-resolution videos in real time. But even with reliable, high-bandwidth communication channels, the combined latency of video encoding, model inference, and round-trip communication prohibits use for certain real-time applications. The alternative is to use fully local inference; but this places extreme constraints on computational and power costs, requiring smaller models and lower resolution, leading to degraded accuracy. To address these challenges, we propose Dedelayed, a real-time inference system that divides computation between a remote model operating on delayed video frames and a local model with access to the current frame. The remote model is trained to make predictions on anticipated future frames, which the local model incorporates into its prediction for the current frame. The local and remote models are jointly optimized with an autoencoder that limits the transmission bitrate required by the available downlink communication channel. We evaluate Dedelayed on the task of real-time streaming video segmentation using the BDD100k driving dataset. For a round trip delay of 100 ms, Dedelayed improves performance by 6.4 mIoU compared to fully local inference and 9.8 mIoU compared to remote inference—an equivalent improvement to using a model ten times larger.
@@ 1 Introduction
Figure 1: Overview of various inference setups, including conventional local, remote, and split inference. Dedelayed combines a small on-device image model and a heavier cloud-based temporally predictive video model to produce accurate and on-time predictions.
In soft real-time applications—such as cloud gaming or video conferencing [jiang2025real, isikdogan2020eye, jain2025ntire]—late outputs may be diminished in value, but are still useful. In these applications, expensive Deep Neural Network (DNN)-based operations can be offloaded to powerful cloud GPUs to save on-device power. As long as the typical latency is low, the loss of utility from latency is outweighed by power savings and extended battery life.
In hard real-time applications—such as aerial robotic control or obstacle avoidance [liu2022large, moon2021real]—late outputs can be catastrophic, and the system must be designed with a guaranteed deadline. Due to the irreducible high-tail latency in wireless communication, hard real-time applications must be equipped with a fully functional local inference pipeline as a fallback in the case that the remote predictions fail to meet the deadline [chen2024fogros2, chen2025fogros2].
Modern hardware for capturing digital video can operate with extremely low power while ingesting super-human amounts of information at high spatial and temporal resolution [engel2023project]. Using video as the primary or exclusive modality of perception provides an opportunity to significantly reduce size, weight, and power by replacing multiple specialized sensors, and is a major focus of ongoing research. Existing video understanding systems based on masked prediction and diffusion-based foundation models [ravi2024sam, agarwal2025cosmos, kim2024openvla] are able to directly produce useful predictions or actions from video input, but require too much computation to run directly on the sensor devices used in these applications. Various approaches to split computing and collaborative inference [kang2017neurosurgeon, matsubara2022split, wang2024end, gao2025feature] have been proposed to offload computation of these expensive image and video models to the cloud. For real-time streaming video applications, existing approaches still fall into one or more of three common pitfalls. (1) They allocate all on-device power and computation to a single linear inference pipeline, leaving no resources for a local-only fallback. (2) They do not account for the impact of latency on prediction accuracy. (3) They operate on videos with significantly reduced spatiotemporal resolution to manage computational cost, leaving out rich visual details available from modern camera systems.
To address these limitations, we introduce Dedelayed (Fig.˜1), a collaborative inference framework suitable for hard real-time applications. Dedelayed is designed to maximally utilize high-resolution, full-motion video understanding models on cloud GPUs, while avoiding the pitfalls of other split-computing systems:
@@
    Temporal prediction for latency mitigation. During supervised training of the remote model component, we simulate a delay of $D$ frames. In other words, the remote model is trained to predict features useful in the future. A delay embedding—similar to a position embedding in text or vision transformers—allows the behavior of the remote model to adapt to changes in the channel. As shown in Fig.˜2, temporally predictive training is able to capture motion dynamics, which can be used to compensate for latency.
@@
    Mixed-resolution inference. On-device AI video processing at capture resolution and frame rate is rarely feasible, even with lightweight models. Dedelayed enables mixed-resolution inference—the local model runs at a lower resolution, and the remote model processes high-resolution frames with a 3D transformer that understands motion. Thus, the remote model supplies delayed yet accurate, high-level semantic features, while the local model aligns and localizes them to the current scene, as shown in Fig.˜3.
@@
    We introduce Dedelayed, a co-inference framework that integrates the output of a future-predicting remote model with the current input to a local model.
@@ 2 Background
Machines equipped with digital video sensors—which are at the center of ongoing innovation in robotics [kim2024openvla, o2024open], remote sensing [szwarcman2024prithvi, khani2021realtimemodelstreaming], and wearable technology [grauman2022ego4d, grauman2024ego]—produce high-bandwidth streams that require significant information processing. The throughput and power efficiency of ingesting pixels on the sensor device (e.g., a battery-powered robot) are extremely high—typically tens or hundreds of megapixels per watt-second [engel2023project]. However, moderately sized DNNs can only process visual data at about one megapixel per watt-second [cai2023efficientvit]. For more advanced video AI based on autoregressive modeling [agarwal2025cosmos] or temporal prediction [assran2025vjepa2], the efficiency may be as low as 500 pixels per watt-second. Instead of on-device processing, power constraints can be circumvented by compressing and transmitting video streams to cloud GPU datacenters supported by a 100-megawatt power infrastructure [goldberg2013cloud].
Nevertheless, fully remote processing is challenging for certain real-time applications (e.g., collision avoidance) due to unreliability in network and cloud infrastructure [chen2024fogros2, chen2025fogros2]. This has motivated split computing systems [teerapittayanon2016branchynet, kang2017neurosurgeon, mpegAI2025, choi2018deepfeaturecompression, choi2018nearlosslessdeepfeaturecompression, choi2022scalableimagecoding, azizian2022privacy] which can reduce the bandwidth and latency of remote inference by leveraging both on-device and remote computation. Still, delivering predictions by a guaranteed deadline requires a fallback procedure independent of the remote server. In many systems (e.g., autonomous motor vehicles) the limited accuracy and reliability of lightweight local models warrant a human operator as the fallback [on2021taxonomy], preventing full automation.
@@ 3 Related work
The challenges of real-time video perception have been approached from many angles. Previous research can be grouped into four major categories: (1) improving the accuracy-efficiency trade-off of inference, (2) offloading computation from the sensor device to a remote server, (3) temporally predictive video modeling, and (4) dynamically augmenting the on-device model via remote collaboration.
Improving the accuracy-efficiency trade-off.
Over the decade since AlexNet [krizhevsky2012alexnet], the trade-off between accuracy and computational efficiency of image-based DNNs has improved significantly. This has made local inference (shown in the first column of Fig.˜1) the de facto standard for many computer vision and robotics applications, since the target perception accuracy is often achievable on-device without introducing the additional failure modes or network requirements. Still, video models that process dense 3D pixel volumes at high resolution using convolutional or transformer-based DNNs are typically avoided for real-time applications. Instead, it is common to apply a lightweight 2D image model to each frame. Using recent 2D models like EfficientViT [cai2023efficientvit], high resolution images can be segmented accurately at real-time throughput (greater than 30 fps) using moderately powered hardware (10–100 Watts). For more restricted power budgets (less than 10 Watts) the spatial or temporal resolution must be reduced by one or more orders of magnitude to meet real-time deadlines, significantly reducing accuracy.
Offloading computation.
Numerous works have explored ways to offload computation of expensive image and video models to the cloud. Early exiting [teerapittayanon2016branchynet], split computing [kang2017neurosurgeon], and feature coding for machines (FCM) [mpegAI2025, choi2018deepfeaturecompression, choi2018nearlosslessdeepfeaturecompression, choi2022scalableimagecoding, azizian2022privacy] focus on partitioning DNN layers into two components: one that runs on the sensor device and another that runs in the cloud. The overall workflow of these approaches is shown in the third column of Fig.˜1. A limitation of these methods and many of their successors is that they allocate local computation towards a prediction pipeline that terminates in the cloud, rather than on the device. This may be desirable for applications where the predictions are only used in the cloud (e.g., remote sensing). However, if predicted outputs must be available on-device (e.g., autonomous navigation), the delay incurred by round-trip communication leads to stale predictions. Even worse, the exclusive allocation of local computational resources to an inference pipeline that necessarily requires a remote server leaves little room for a readily available local fallback mode in the event of late or dropped predicted outputs.
Temporally predictive video modeling.
Regardless of DNN size or added network latency, some amount of inference delay is unavoidable. Instead of focusing exclusively on reducing the degree of delay, prior works have explored correcting the estimated world state based on the anticipated delay. These corrections require some mechanism for temporally predictive video modeling, and can use on-device optical flow estimates [zhang2024neuflow] or readily available motion vectors from an existing video codec [xiao2024task]. If video is streamed to a sufficiently powerful remote server, generative video world models based on masked prediction [assran2025vjepa2] or diffusion [agarwal2025cosmos] foundation models can provide temporal prediction. In each of these cases, existing methods for temporal prediction are repurposed from another application, rather than specialized and optimized for the desired perception task. This prevents temporally predictive abilities or behaviors from being learned or optimized with the rest of the DNN-based layers.
Dynamically augmenting the on-device model via remote collaboration during streaming inference.
This category of methods improve the on-device model’s inference capabilities by incorporating remote information at runtime (e.g., streaming weights or fusing remote features). They are underexplored but important for real-time applications on computationally limited devices. Adaptive Model Streaming [khani2021realtimemodelstreaming] streams weight updates to keep the local model tuned to changing scene characteristics. Knowledge Boosting [srinivas2024knowledgeboosting] addresses the problem from a different lens—instead of dynamically updating the model weights, it directly fuses in remote features into a small on-device model. It requires separate models trained for specific fixed delays, and evaluates on audio data in a controlled local environment with delays up to 48 ms within a local networking context. Our method, Dedelayed, is capable of adapting to arbitrary remote inference delays by conditioning the remote model on the measured delay at runtime, enabling dynamic specialization across varying latency. Methods within this category are orthogonal to and compatible with many methods from earlier categories, and with each other. Our work extends this rarely-explored direction by explicitly inputting dynamic system characteristics such as delay to improve on-device predictions under realistic, variable latency.
@@ 4 Method
Dedelayed introduces a general framework that improves the accuracy and robustness of real-time inference on resource-constrained sensor devices. It does so by combining the strengths of both local inference and remote inference, while mitigating their weaknesses. The local model has access to the latest sensor data, and yet lacks the computational capability needed to produce accurate outputs. The remote model provides accurate outputs, and yet delivers them with delay. By careful combination of both subsystems, Dedelayed is able to provide bounded performance guarantees. When trained appropriately such that the local and remote subsystems deliver the maximum accuracies that they would conventionally—e.g., by freezing the local and remote image backbones during the final joint training of the entire system—it can be guaranteed that Dedelayed is effectively never worse than either local inference or remote inference independently. As we will demonstrate later, we are able to glue together the two subsystems in a way that is simple yet effective.
@@
Dedelayed can be formulated in simple mathematical terms as follows. Given a fresh input frame $x_{t}$ at current time $t$, the final prediction $\hat{y}_{t}$ is computed using a lightweight local model, $f_{\text{local}}$, which processes $x_{t}$ along with time-delayed features $z_{t-\tau}$ from a heavyweight remote model, $f_{\text{remote}}$. To produce powerful predictive features, the remote model is conditioned on the delay $\tau$, and processes a short clip of past frames $x_{\leq t-\tau}$ ending at time $t-\tau$. This is expressed by the following equations:
$z_{t-\tau}=f_{\text{remote}}(\tau,\,x_{\leq t-\tau})$ (1)
$\hat{y}_{t}=f_{\text{local}}(x_{t},\,z_{t-\tau})$ (2)
For clarity, the notation is summarized in Tab.˜1. Fig.˜4 presents a system diagram that demonstrates the fundamental principle we describe, and shows how information propagates through the various subsystems as time progresses.
@@
$z_{t-\tau}$ Features outputted by remote model run at time $t-\tau$
$\tau$ Delay in time between the old and current frame
$D$ Delay in frames between the old and current frame
$f_{\text{local}}$ Light local model run at time $t$
$f_{\text{remote}}$ Heavy remote model run at time $t-\tau$
@@
Figure 5: Dedelayed workflow. The complementary properties of local and remote inference are exploited to produce real-time streaming video understanding that is both accurate and on time.
5 Design and implementation
Dedelayed aims to use remote computation to improve the accuracy and robustness of real-time machine vision on resource constrained sensor devices. The design of Dedelayed is guided by three key insights. (1) Incorporating the outputs of the remote model as side information to the local model eliminates computational redundancy and provides a fallback suitable for hard real-time applications. (2) Temporally predictive training can mitigate the impact of latency due to processing and round-trip communication. (3) Decoupling inference into a high resolution full-motion video stream and a low-resolution still-image stream exploits the complementary properties of remote and on-device processing. Fig.˜5 shows an overview of the design in the context of semantic segmentation of video frames.
@@ 5 Design and implementation
    Each incoming frame is fed into a heavyweight remote model $f_{\text{remote}}$. To accelerate inference, this model caches and maintains a context window of the $K$ most recent features, which are computed from each incoming frame using a pretrained 2D ViT backbone.
@@
    The lightweight local model $f_{\text{local}}$ runs on a fresh input $x_{t}$, and fuses in the remote features $z_{t-\tau}$.
@@
Local image model incorporating remote side information
The local model processes the most recently collected video frame on the sensor device at low resolution, and consists of three learnable modules, shown in the right half of Fig.˜5: (1) pixel-processing 2D convolutional layers (CNN2D), (2) 2D feature-processing convolution and attention layers (CoAt2D), and task specific (e.g., classification or segmentation) MLP layers (L-MLP). If present, features from the remote model are added element-wise to the activation map between the CNN2D and the CoAt2D. Our local model can operate at a resolution up to 0.34 megapixels ($704\times 480$) while meeting the target latency of 33 ms (a single frame at 30 fps) on our CPU testbed.
Video compression
Each frame captured on the sensor device is processed by a lossy image or video codec to allow transmission over a wireless channel. In our experiments, we choose the resolution, framerate, and degree of lossy compression to represent a video signal that is transmissible via 5G cellular uplink (30fps 720p content compressed at rates between 1–10 Mbps).
Latency-aware remote video model
The remote model operates on high resolution, compressed video frames, and consists of four learnable modules: (1) a 2D vision transformer (ViT2D), (2) a 3D video transformer (ViT3D), (3) task-specific MLP layers (R-MLP), and (4) a dimensionality-reducing autoencoder (DR-AE). During the initial training stages, only modules 1–3 (ViT2D, ViT3D, and R-MLP) are used, as shown in Fig.˜6. During the final training stage, the R-MLP module is discarded and replaced with the DR-AE to allow joint training of the local and remote components. Due to round trip communication and networking, the predictions from the remote model will be delayed by $D$ frames at runtime. During training, an artificial delay is applied to the input, but not to the target. Thus, the training objective is to predict the target of a future frame. During training, we sample a delay uniformly between 0 and 5 frames (up to 167 ms). Additionally, we add a learnable “delay embedding” to the input activation maps of the ViT3D—similar to the position or timestep embedding used in diffusion transformers [peebles2023scalable]. In our experiments, we provide the remote model with four frames of context and aim for the same target latency (33 ms), but on a high power GPU (rather than CPU) testbed.
Figure 6: Overview of the remote model component as used during pre-training. A video sequence $v[n]$ is captured on the sensor device and transmitted to a remote server, incurring a delay of $D$ frames. The remote model operates on a fixed context window of $K$ frames. A 2D vision transformer (EfficientViT-L1, effective patch size of $8\times 8$ pixels) is applied independently to each of the $K$ input frames. The outputs of the 2D transformer are concatenated along the temporal axis, but spatially merged into larger $16\times 16$ patches, thus maintaining a similar sequence length (when $K=4$, the sequence length is identical). A learned delay embedding is added, allowing the remaining layers to behave differently based on the expected value of $D$. Finally, 3D video transformer layers and task-specific (e.g., segmentation) layers are applied, generating a prediction corresponding to the current frame ($D$ frames in the future after the latest remote input frame).
Joint prediction with autoencoded resolution adapter.
After pretraining the remote model for the desired prediction task, its task-specific layers are discarded and replaced with the DR-AE, consisting of adaptive spatial pooling and a channel bottleneck. The degree of adaptive spatial pooling is chosen to match the operating resolution of the local model. The sample mean is used as the pooling operator. The remainder of the DR-AE consists of MLP layers that terminate in a low-resolution, low-channel count activation map suitable for downlink transmission. The output of the DR-AE is added element-wise to the output of the pixel-processing CNN2D. Then, the local and remote models undergo a final joint training stage.
6 Evaluation